The lifetime T (years) of an electronic component is a continuous random variable with a


probability density function given by


( ) , 0 t


f t e t 


 


(a) Find the lifetime L which a typical component is 60% certain to exceed.


(b) If five components are sold to a manufacturer, find the probability that at least one of


them will have a lifetime less than L years

1
Expert's answer
2021-11-01T13:05:19-0400

The lifetime "T" (years) of an electronic component is a continuous random variable with a probability density function given by

"f(t)=e^{-t}, t\geq0"

(a)


"P(T>L)=\displaystyle\int_{L}^{\infin}e^{-t}dt=\lim\limits_{A\to \infin}\displaystyle\int_{L}^{A}e^{-t}dt"

"=\lim\limits_{A\to \infin}\displaystyle[-e^{-t}]\begin{matrix}
 A \\
 L
\end{matrix}=e^{-L}"

Given "P(T>L)=0.6." Then


"e^{-L}=0.6"

"L=-\ln0.6"

"L\approx0.511\ years."

(b) Assuming that the lifetime of each component is independent we have


"P(at\ least\ one\ component\ has \ T<0.511)"

"=1-P(no\ component\ has \ T<0.511)"

"=1-(0.6)^5=0.92224"